OpenStudy (anonymous):

with the x-intercept as 5x-6y=-15 parallel to 2x-8y=9

OpenStudy (anonymous):

Is part of the question missing? What comes before the "with"?

OpenStudy (anonymous):

its solving equations of a line and putting it in standard form

OpenStudy (anonymous):

In the mean time, 2x - 8y = 9 - 8y = 9 - 2x y = -9/8 + x/4

OpenStudy (anonymous):

OK, so we already have one standard form from above. x/4 = (1/4)x so the slope of that particular line is 1/4. The parallel line will have the same slope. To find x-intercept, we plug 0 in for y right?

OpenStudy (anonymous):

howd you get x/4?

OpenStudy (anonymous):

(-2x)/(-8) right?

OpenStudy (anonymous):

5x - 6(0) = -15 ==> x =? When you have this value then you can say: because this is the x-intercept of the line, the line goes through the point (?, 0) Now you have the classic problem of finding the line given slope and 1 point. You can do that right?

OpenStudy (anonymous):

First. we determine the x-intercept of 5x - 6y = -15 (a,0) when y = 0, x = -3, so x intercept is (-3, 0) we use the form y = mx + b since the line is parallel to 2x - 8y = 9, it will have the same slope rewriting the equation in slope intercept form y = (2/8)x - (9/8) slope is 2/8 equation now is y = (2/8)x + b we now need to find b, which is the y-intercept using that point (-3,0), we can plug that in y = (2/8)x + b and can now solve for b 0 = (2/8)(-3) + b I guess you can do the rest.